OpenStudy (anonymous):

Solve by taking square roots -8+36r^2=1

thomaster (thomaster):

first you'd have to get -8 to the right

OpenStudy (anonymous):

so add 8 to both sides.. i got 36r^2=9

thomaster (thomaster):

Yes, any idea what to do next?

OpenStudy (anonymous):

divide 36

thomaster (thomaster):

yea divide 9 by 36

OpenStudy (anonymous):

.25?

thomaster (thomaster):

Yep. So now we have \(\Large r^2=\frac{1}4\)

thomaster (thomaster):

The last step would be take the square root of both sides to solve for r

OpenStudy (anonymous):

and how would i get an answer to that?

thomaster (thomaster):

\(\Large\sqrt{0.25}\) Use a calculator

OpenStudy (anonymous):

.5

thomaster (thomaster):

yep :) So \(\Large r=\frac{1}2\) (or 0.5)

OpenStudy (anonymous):

Goodness, thank you so much! :D

thomaster (thomaster):

haha no problemo ◠◡◠